1)how to solve the equation x^2+4x-77=0
2) factor completely 27x-8x^4
3) factor completely 64x^2+112xy+49y^2

Answer by checkley77(12844)

1)how to solve the equation
x^2+4x-77=0
(X-7)(X+11)=0
X-7=0
X=7 ANS
X+11=0
X=-11 ANS.
-----------------------------
2) factor completely
27x-8x^4
X(3-2X)(9+6X+4X^2) ANS.
3) factor completely
64x^2+112xy+49y^2
(8X+7Y)^2 ANS.